# DreamPhysics: Learning Physics-Based 3D Dynamics with Video Diffusion Priors

### v2

@@ Abstract
Dynamic 3D interaction has been attracting a lot of attention recently. However, creating such 4D content remains challenging. One solution is to animate 3D scenes with physics-based simulation, which requires manually assigning precise physical properties to the object or the simulated results would become unnatural. Another solution is to learn the deformation of 3D objects with the distillation of video generative models, which, however, tends to produce 3D videos with small and discontinuous motions due to the inappropriate extraction and application of physical prior. In this work, combining the strengths and complementing shortcomings of the above two solutions, we propose to learn the physical properties of a material field with video diffusion priors, and then utilize a physics-based Material-Point-Method (MPM) simulator to generate 4D content with realistic motions. In particular, we propose motion distillation sampling to emphasize video motion information during distillation. Moreover, to facilitate the optimization, we further propose a KAN-based material field with frame boosting. Experimental results demonstrate that our method enjoys more realistic motion than state-of-the-arts. Codes are released at: https://github.com/tyhuang0428/DreamPhysics.
Figure 1: (a): The values of physical properties can significantly affect the quality of the simulated videos. (b) Using state-of-the-art video diffusion models (Blattmann et al. 2023; Wang et al. 2023c) can hardly generate the desired results. (c) Our DreamPhysics can produce realistic 3D dynamic content with the distillation of video diffusion priors.
@@ Introduction
With the development in 3D representation, e.g., Neural Radiance Fields (NeRF) (Mildenhall et al. 2021) and 3D Gaussian Splatting (GS) (Kerbl et al. 2023), significant progress has been made in creating 3D assets through reconstruction and generation (Poole et al. 2022; Wang et al. 2024). However, interacting with these 3D assets in a simulation environment (Savva et al. 2019; Xia et al. 2018) remains challenging, while it is critical in various applications, e.g., video games (Fan et al. 2022), virtual reality (Jiang et al. 2024), and robotic manipulation (Lu et al. 2024).
@@
Animating static 3D objects based on instructions is an important step toward this interaction goal. In the physical world, object movement is intertwined with the object’s internal properties (e.g., material types). Hence, we can see that on the one hand, some works (Xie et al. 2023; Feng et al. 2024b) first inject physical parameters into 3D GS objects, and then perform motion predictions in a physics-based simulator. However, as all these parameters have to be manually assigned, it is difficult to set them accurately, thus producing unnatural simulation results, as demonstrated in Figure 1(a). On the other hand, pre-trained video generators (Singer et al. 2022; Khachatryan et al. 2023; Wang et al. 2023b) have been trained with real-world video data that naturally considers physical phenomena and regulations. These generators should contain, to some extent, physics-based prior knowledge. Thus, some works (Singer et al. 2023; Bahmani et al. 2023; Zhao et al. 2023b) directly learn time-dependent deformation with the distillation of video diffusion models. However, the generated motions tend to present small and discontinuous motions across frames. We hypothesize that the main reason for this is the inappropriate extraction and application of the physical prior, rather than the utilization of video models. We therefore ask this question: how can we mine and apply the physical knowledge of the video generative models to achieve realistic dynamic 3D synthesis?
@@
To this end, we rethink the usage of physics-based simulation and video generative models in this work. We propose to learn a material field, rather than a deformation field, from video diffusion models, and then deploy a physics-based simulator to animate the 3D object in this field. As such, the advantages of the above two related approaches are combined, while their shortcomings can be complemented. Learnable physical properties from video diffusion models eliminate the need for manual modulation, and the physical simulator based on reasonable properties ensures more realistic motion generation.
@@
Specifically, we introduce a new framework named DreamPhysics. DreamPhysics takes 3D GS (Kerbl et al. 2023) as a 3D representation. It first learns the physical properties of a material field with the distillation of video diffusion priors, and then adopts a simulator based on Material Point Method (MPM) (Stomakhin et al. 2013; Jiang et al. 2016) to model the time-dependent deformation of each Gaussian kernel. During distillation from video diffusion models, common Score Distillation Sampling (SDS) (Poole et al. 2022) may focus more on color information and is not completely suitable for extracting motion information. Instead, we propose motion distillation sampling (MDS) to avoid the interference of color bias and emphasize the motion information in the rendered video. In addition, directly optimizing the material field can easily lead to unstable training due to the large range of possible parameter values. To facilitate the training process, we propose a KAN-based (Liu et al. 2024) material field with frame boosting.
@@
  • We introduce a physics-based 3D animation framework, i.e., DreamPhysics, which learns a material field for a physical simulator to support the creation of dynamic 3D content.
@@
3D supervised methods (Nichol et al. 2022; Cheng et al. 2023; Wei et al. 2023; Jun and Nichol 2023; Yu et al. 2023; Huang et al. 2023b; Hong et al. 2023; Tang et al. 2024) utilize text-3D data to train generators capable of directly producing 3D assets. For instance, Point-E (Nichol et al. 2022) is an early example of a text-to-3D generator that creates point clouds based on input prompts. Shap-E (Jun and Nichol 2023) and LGM (Tang et al. 2024) have expanded the scope of generated content to include SDF (Park et al. 2019) and 3DGS (Kerbl et al. 2023) representations, respectively. Despite their efficiency in generating solid 3D content, these methods are significantly limited by the availability of 3D data. The current scale of 3D training datasets (Reizenstein et al. 2021; Deitke et al. 2023, 2024) is much smaller compared to 2D or video datasets, resulting in a constrained open-world capability relative to image or video generators. TextField3D (Huang et al. 2023b) attempts to enhance text control in 3D generators using a noisy latent space, yet it still falls short of achieving the imaginative capabilities seen in 2D generators.
@@
To address this issue, recent works (Liu et al. 2023a; Zhao et al. 2023a; Shi et al. 2023; Liu et al. 2023b; Long et al. 2023) have explored the synthesis of multi-view images of 3D objects. For example, Zero-1-to-3 (Liu et al. 2023a) generates images of the same object from different viewpoints based on a given image and viewpoint angles. MVDream (Shi et al. 2023) enhances consistency by generating orthogonal multi-view images of the same object. Wonder3D (Long et al. 2023) supports depth generation to achieve more precise object reconstruction.
@@
As the success of video generative models, some methods (Zhao et al. 2023b) attempt to leverage video diffusion models to guide the prediction of the 3D deformation. DreamGuassian4D (Ren et al. 2023) uses a pre-generated video to supervise the deformation of static scenes. Animate124 (Zhao et al. 2023b) proposes to distill the priors of video diffusion models to its deformation fields.
@@
Point cloud (Guo et al. 2020) is an explicit 3D representation, which generally consists of the coordinates for all points. Normal and color information (Dai et al. 2017; Qi et al. 2017) can also be considered to further enrich the feature space of point cloud. Despite the succinct representation, its rendering quality is heavily restricted by the number of points (Zhang et al. 2022; Huang et al. 2023a). Derived from NeRF (Mildenhall et al. 2021), 3D Gaussian Splatting (GS) (Kerbl et al. 2023) introduced a point-based explicit radiance field. Points are modeled as a set of Gaussian kernels $\{\mathcal{G}_{i}\}=\{x_{i},\sigma_{i},\Sigma_{i},C_{i}\}$, where $x_{i}$, $\sigma_{i}$, $\Sigma_{i}$, and $C_{i}$ denote the center coordinate, opacity, covariance matrix, and spherical harmonic coefficient of the $i$-th kernel $\mathcal{G}_{i}$. To render a 3D GS scene at a specific viewpoint $\mathbf{r}$, the color can be formulated as,
@@
The material point method (MPM) (Zong et al. 2023) is a numerical simulation mechanic for the analysis of continuum forces. In MPM, the continuum is represented by a set of particles placed in a grid-based space. Different from mesh-based numerical mechanics, MPM can be naturally applied to point-based representation 3D GS. Following PhysGaussian (Xie et al. 2023), we have a time-dependent state for each Gaussian kernel as:
@@
Figure 2: Overview of DreamPhysics. First, a set of physical parameters is initialized with a KAN-based (Liu et al. 2024) material field for a static 3D GS (Kerbl et al. 2023). Then, it is fed to an MPM simulator to render a 3D video. Finally, we leverage motion distillation sampling to optimize the rendered video, and the distillation gradients are back-propagated to refine the physical parameters.
@@
The score distillation sampling (SDS) (Poole et al. 2022; Wang et al. 2023a) distills pre-trained 2D diffusion models to the parameters of 3D representation, widely used in 3D generation methods. Recently, SDS has had various extensions. Variational score Distillation (VSD) (Wang et al. 2024) proposes an additional LoRA term $\bm{\epsilon}_{\theta}$ to learn the distribution of current 3D scenes, which is attached to the score as,
@@
where $t$ is the noise timestep and $y$ is the input condition. $\hat{\bm{\epsilon}}_{\text{2D}}$ and $\hat{\bm{\epsilon}}_{\theta}$ are noises predicted by a pre-trained 2D diffusion model and the LoRA. Another extension, SDS-T, is for dynamic 3D generation, where video diffusion models are deployed to supervise the time-dependent deformation of static 3D objects. Specifically, given a camera trajectory $\mathbf{r}(t)$, SDS-T optimizes the rendered 3D video $V_{\mathbf{r}(t)}$ with predicted noise $\hat{\epsilon}_{\text{V}}$,
@@
Video generative models are trained with real-world captured videos that cover kinds of physical phenomena. As a result, given a simulated video $V$, we can assess whether it adheres to physical laws based on the judgement of video models. To this end, one direct solution is to treat videos generated by video models as ground truth, supervising $V$ with reconstruction loss (Zhang et al. 2024). However, limited by the control capability, existing video generators can hardly produce desired ground-truth videos. We consider exploring distillation methods to optimize simulated results. Motion distillation sample is thus proposed to enhance the distillation of video diffusion’s motion priors.
@@
Note that the gradient of $s_{\text{MDS}}$ cannot be directly propagated to the target physical parameters $\theta_{\mathcal{G}}$, and it needs to go through the differentiable MPM. Thus, our training objective can be written as,
@@
The value range for physical properties $\theta_{\mathcal{G}}$ can be very large, e.g., the reasonable values for Young’s modulus can vary from $1$ to $71e$ 7. However, during gradient updates, the same gradient can result in varying update granularity across different magnitudes, causing parameters to get stuck within a specific magnitude range. To enable parameters to converge more quickly to a reasonable range, we propose to perform a KAN-based tri-plane representation to model the material field and conduct frame boosting to further facilitate the training process.
KAN-Based Triplane. Tri-plane is widely used to encode spatial information. Given a 3D coordinate $x$, the tri-plane extractor projects it onto three orthogonal planes, i.e., the front view, side view, and top view. These projections match $x$ with 2D features that represent different perspectives of the 3D space. We extract features with KAN (Liu et al. 2024), which integrates kernel methods and attention mechanisms to offer superior modeling capabilities for physics-based tasks compared to traditional MLPs. Extracted features are then combined to form a unified representation, which constitutes our physical parameters $\theta_{\mathcal{G}}$. The gradient is propagated as,
@@
Frame Boosting. The MPM simulator is a sequential model, which can easily lead to gradient vanishing or exploding like RNN (Rumelhart, Hinton, and Williams 1986). We have to conduct truncated back-propagation through time (BPTT), preserving the gradient of key frame simulation only. Truncated BPTT can effectively prevent gradient issues, but the supervision could be limited to specific frames. To ensure that our supervision covers as many video frames as possible, we further suggest a frame-boosting strategy. Specifically, given a total number of frames $M\times T$, we can separate them into $M$ groups of frames with equal intervals, i.e., $V_{t_{i}}=\{I_{i},I_{i+M},...,I_{i+M(T-1)}\}$ for the $i$-th group. These groups formulate different videos, which are fed into the supervision process alternately. Finally, the boosted motion distillation can be formulated as,
@@
Implementation Details. The simulation is based on the warp (Macklin 2022) implementation of MPM (Zong et al. 2023). For most simulation scenes, we set the simulation duration as $5\times 10^{-5}$ second and the frame duration as $4\times 10^{-2}$ second. Thus, we simulate 800 steps between every two renderings and include the simulation gradient of the last step in the optimization. We leverage a text-to-video diffusion model ModelScope (Wang et al. 2023b) and an image-to-video diffusion model Stable Video Diffusion (SVD) (Blattmann et al. 2023) to conduct text-conditioned and image-conditioned optimization, respectively. The numbers of their generated video frames $T$ are 16 and 25, respectively. For frame boosting, we set $M=5$, boosting the video slices to 5 groups. The setting of MDS follows SDS, where CFG value is set to 100. We stop the training if optimized parameter values stabilize within one order of magnitude. The training process requires around 30 iterations. The iteration time highly depends on the number of input Gaussian kernels, and it is within 30 seconds for most cases.
@@
Figure 3: (a) Text-conditioned optimization. The ficus will excessively tilt to one side before optimization; (b) Image-conditioned optimization. Right images are the space-time (X-t) slices, one axis represents time and the other axis shows a space slice (red line) of the object.
@@
Comparison with State-of-the-art Works. We report the quantitative results of all the compared methods in Table 1. Since PhysDreamer hasn’t released its training implementation, we can only compare four evaluation scenes, which are provided in PhysDreamer’s video demo. Considering that other methods don’t have extra text inputs, we use the first frame as the image condition to conduct the optimization. According to the evaluation of aesthetic quality, our results are the closest to the ground truth. PhysDreamer has a lower score compared with PhysGaussian, which indicates that pre-generated videos may not be a proper ground truth for supervision. The generation quality of DreamGaussian4D is the worst because its deformation prediction didn’t consider physical constraints.

### v3

@@ Abstract
Dynamic 3D interaction has been attracting a lot of attention recently. However, creating such 4D content remains challenging. One solution is to animate 3D scenes with physics-based simulation, which requires manually assigning precise physical properties to the object or the simulated results would become unnatural. Another solution is to learn the deformation of 3D objects with the distillation of video generative models, which, however, tends to produce 3D videos with small and discontinuous motions due to the inappropriate extraction and application of physics priors. In this work, to combine the strengths and complementing shortcomings of the above two solutions, we propose to learn the physical properties of a material field with video diffusion priors, and then utilize a physics-based Material-Point-Method (MPM) simulator to generate 4D content with realistic motions. In particular, we propose motion distillation sampling to emphasize video motion information during distillation. In addition, to facilitate the optimization, we further propose a KAN-based material field with frame boosting. Experimental results demonstrate that our method enjoys more realistic motions than state-of-the-arts do.
Figure 1: (a): The setting of physical properties can significantly affect the quality of the simulated videos. (b) Using state-of-the-art video diffusion models (Blattmann et al. 2023; Wang et al. 2023c) can hardly generate the desired results. (c) Our DreamPhysics can produce realistic 3D dynamic content with the distillation of video diffusion priors.
Codehttps://github.com/tyhuang0428/DreamPhysics
@@ Introduction
With the development in 3D representations, e.g., Neural Radiance Fields (NeRF) (Mildenhall et al. 2021) and 3D Gaussian Splatting (GS) (Kerbl et al. 2023), significant progress has been made in creating 3D assets through reconstruction and generation (Poole et al. 2022; Wang et al. 2024). However, interacting with these 3D assets in a simulation environment (Savva et al. 2019; Xia et al. 2018) remains challenging, despite its importance in many applications, e.g., video games (Fan et al. 2022), virtual reality (Jiang et al. 2024), and robotics (Lu et al. 2024).
@@
Animating static 3D objects based on instructions is an important step toward this interaction goal. In the real world, object movement is intertwined with the object’s internal properties (e.g., material types). Hence, we can see that on the one hand, some works (Xie et al. 2023; Feng et al. 2024b) first inject physical parameters into 3D GS objects, and then perform motion predictions in a physics-based simulator. However, as all these parameters have to be manually assigned, it is difficult to set them accurately, thus producing unnatural simulation results, as demonstrated in Figure 1(a). On the other hand, pre-trained video generators (Singer et al. 2022; Khachatryan et al. 2023; Wang et al. 2023b) are trained on real-world video data, which has naturally incorporated physical phenomena and regulations. These generators should contain, to some extent, physics-based prior knowledge. Thus, some works (Singer et al. 2023; Bahmani et al. 2023; Zhao et al. 2023) directly learn time-dependent deformation with the distillation of video diffusion models. However, the generated motions tend to exhibit small and discontinuous motions across frames. We hypothesize that the main reason for this drawback is the inappropriate extraction and application of the physics prior, rather than the utilization of video models. We therefore ask this question: how can we mine and apply the physics knowledge of video generative models to achieve realistic dynamic 3D synthesis?
@@
To this end, we rethink the usage of physics-based simulation and video generative models in this work. We propose to learn a material field, rather than a deformation field, from video diffusion models, and then deploy a physics-based simulator to animate the 3D object in this field. As such, the advantages of the above two related approaches are combined, while their shortcomings can be complemented. Learnable physical properties from video diffusion models eliminate the need for manual modulation, and the physics simulator based on reasonable properties ensures more realistic motion generation.
@@
Specifically, we introduce a new framework named DreamPhysics. DreamPhysics takes 3D GS (Kerbl et al. 2023) as a 3D representation. It first learns the physical properties of a material field with the distillation of video diffusion priors, and then adopts a simulator based on Material Point Method (MPM) (Stomakhin et al. 2013; Jiang et al. 2016) to model the time-dependent deformation of each Gaussian kernel. During the distillation from video diffusion models, the Score Distillation Sampling (SDS) (Poole et al. 2022) may focus more on color information, and is not completely suitable for extracting motion information. Instead, we propose motion distillation sampling (MDS) to avoid the interference of color bias and emphasize the motion information in the rendered video. In addition, directly optimizing the material field can easily lead to unstable training due to the large range of possible parameter values. To facilitate the training process, we propose a KAN-based (Liu et al. 2024) material field with frame boosting.
@@
  • We introduce a physics-based 3D animation framework, i.e., DreamPhysics, which learns a material field for a physics simulator to support the creation of dynamic 3D content.
@@
3D supervised methods (Nichol et al. 2022; Jun and Nichol 2023; Yu et al. 2023; Huang et al. 2023b; Hong et al. 2023; Tang et al. 2024) utilize text-3D data to train generators capable of directly producing 3D assets. For instance, Point-E (Nichol et al. 2022) is an early example of a text-to-3D generator that creates point clouds based on input prompts. Shap-E (Jun and Nichol 2023) and LGM (Tang et al. 2024) have expanded the scope of generated content to include SDF (Park et al. 2019) and 3DGS (Kerbl et al. 2023) representations, respectively. Despite their efficiency in generating solid 3D content, these methods are significantly limited by the availability of 3D data. The current scale of 3D training datasets (Reizenstein et al. 2021; Deitke et al. 2023, 2024) is much smaller compared to 2D or video datasets, resulting in a constrained open-world capability relative to image or video generators. TextField3D (Huang et al. 2023b) attempts to enhance text control in 3D generators using a noisy latent space, yet it still falls short of achieving the imaginative capabilities seen in 2D generators.
@@
To address this issue, recent works (Liu et al. 2023; Shi et al. 2023; Long et al. 2023) have explored the synthesis of multi-view images of 3D objects. For example, Zero-1-to-3 (Liu et al. 2023) generates images of the same object from different viewpoints based on a given image and viewpoint angles. MVDream (Shi et al. 2023) enhances consistency by generating orthogonal multi-view images of the same object. Wonder3D (Long et al. 2023) supports depth generation to achieve more precise object reconstruction.
@@
As the success of video generative models (Wang et al. 2023c, b; Blattmann et al. 2023; Zhang et al. 2023), some methods (Zhao et al. 2023) attempt to leverage video diffusion models to guide the prediction of the 3D deformation. DreamGuassian4D (Ren et al. 2023) uses a pre-generated video to supervise the deformation of static scenes. Animate124 (Zhao et al. 2023) proposes to distill the priors of video diffusion models to its deformation fields.
@@
Point cloud (Guo et al. 2020) is an explicit 3D representation, which generally consists of the coordinates for all points. Normal and color information (Dai et al. 2017; Qi et al. 2017) can also be considered to further enrich the feature space of the point cloud. Despite the succinct representation, its rendering quality is heavily restricted by the number of points (Huang et al. 2023a). Derived from NeRF (Mildenhall et al. 2021), 3D Gaussian Splatting (GS) (Kerbl et al. 2023) introduces a point-based explicit radiance field. Points are modeled as a set of Gaussian kernels $\{\mathcal{G}_{i}\}=\{x_{i},\sigma_{i},\Sigma_{i},C_{i}\}$, where $x_{i}$, $\sigma_{i}$, $\Sigma_{i}$, and $C_{i}$ denote the center coordinate, opacity, covariance matrix, and spherical harmonic coefficient of the $i$-th kernel $\mathcal{G}_{i}$. To render a 3D GS scene at a specific viewpoint $\mathbf{r}$, the color can be formulated as:
@@
The material point method (MPM) (Stomakhin et al. 2013; Jiang et al. 2016) is a numerical simulation mechanic for the analysis of continuum forces. In MPM, the continuum is represented by a set of particles placed in a grid-based space. Different from mesh-based numerical mechanics, MPM can be naturally applied to point-based representation 3D GS. Following PhysGaussian (Xie et al. 2023), we have a time-dependent state for each Gaussian kernel as:
@@
Figure 2: Overview of DreamPhysics. First, a set of physical parameters is initialized with a KAN-based material field for a static 3D GS. Then, it is fed to an MPM simulator to render a 3D video. Finally, we leverage motion distillation sampling to optimize the rendered video, and the distillation gradients are back-propagated to refine the physical parameters.
@@
The score distillation sampling (SDS) (Poole et al. 2022; Wang et al. 2023a) distills pre-trained 2D diffusion models to the parameters of the 3D representation, widely used in 3D generation methods. Recently, SDS has had various extensions. Variational score Distillation (VSD) (Wang et al. 2024) proposes an additional LoRA term $\bm{\epsilon}_{\theta}$ to learn the distribution of current 3D scenes, which is attached to the score as:
@@
where $t$ is the noise timestep and $y$ is the input condition. $\hat{\bm{\epsilon}}_{\text{2D}}$ and $\hat{\bm{\epsilon}}_{\theta}$ are noises predicted by a pre-trained 2D diffusion model and the LoRA. Another extension, SDS-T, is for dynamic 3D generation, where video diffusion models are deployed to supervise the time-dependent deformation of static 3D objects. Specifically, given a camera trajectory $\mathbf{r}(t)$, SDS-T optimizes the rendered 3D video $V_{\mathbf{r}(t)}$ with predicted noise $\hat{\epsilon}_{\text{V}}$, as:
@@
Video generative models are trained with real-world captured videos that cover kinds of physical phenomena. As a result, given a simulated video V, we can assess whether it is natural and realistic based on the judgement of video models. To this end, one direct solution is to treat videos generated by video models as ground truth, supervising $V$ with reconstruction loss (Zhang et al. 2024). However, limited by the control capability, existing video generators can hardly produce desired ground-truth videos. We consider exploring distillation methods to optimize simulated results. Motion distillation sample is thus proposed to enhance the distillation of video diffusion’s motion priors.
@@
Note that the gradient of $s_{\text{MDS}}$ cannot be directly propagated to the target physical parameters $\theta_{\mathcal{G}}$, and it needs to go through the differentiable MPM. Thus, our training objective can be written as:
@@
The value range for physical properties $\theta_{\mathcal{G}}$ can be very large, e.g., the reasonable values for Young’s modulus can vary from $41e$ 4 to $81e$ 8. However, during gradient updates, the same gradient can result in varying update granularity across different magnitudes, causing parameters to get stuck within a specific magnitude range. To enable parameters to converge more quickly to a reasonable range, we propose to perform a KAN-based tri-plane representation to model the material field and conduct frame boosting to further facilitate the training process.
KAN-Based Triplane. Tri-plane is widely used to encode spatial information. Given a 3D coordinate $x$, the tri-plane extractor projects it onto three orthogonal planes, i.e., the front view, side view, and top view. These projections match $x$ with 2D features that represent different perspectives of the 3D space. We extract features with KAN (Liu et al. 2024), which integrates kernel methods and attention mechanisms to offer superior modeling capabilities for physics-based tasks compared to traditional MLPs. Extracted features are then combined to form a unified representation, which constitutes our physical parameters $\theta_{\mathcal{G}}$. The gradient is propagated as:
@@
Frame Boosting. The MPM simulator is a sequential model, which can easily lead to gradient vanishing or exploding like RNN (Rumelhart, Hinton, and Williams 1986). We have to conduct truncated back-propagation through time (BPTT), preserving the gradient of key frame simulation only. Truncated BPTT can effectively prevent gradient issues, but the supervision could be limited to specific frames. To ensure that our supervision covers as many video frames as possible, we further suggest a frame-boosting strategy. Specifically, given a total number of frames $M\times T$, we can separate them into $M$ groups of frames with equal intervals, i.e., $V_{t_{i}}=\{I_{i},I_{i+M},...,I_{i+M(T-1)}\}$ for the $i$-th group. These groups formulate different videos, which are fed into the supervision process alternately. Finally, the boosted motion distillation can be formulated as:
@@
Implementation Details. The simulation is based on the warp (Macklin 2022) implementation of MPM (Stomakhin et al. 2013; Jiang et al. 2016). For most simulation scenes, we set the simulation duration as $5\times 10^{-5}$ second and the frame duration as $4\times 10^{-2}$ second. Thus, we simulate 800 steps between every two renderings and include the simulation gradient of the last step in the optimization. We leverage a text-to-video diffusion model ModelScope (Wang et al. 2023b) and an image-to-video diffusion model Stable Video Diffusion (SVD) (Blattmann et al. 2023) to conduct text-conditioned and image-conditioned optimization, respectively. The numbers of their generated video frames $T$ are 16 and 25, respectively. For frame boosting, we set $M=5$, boosting the video slices to 5 groups. The setting of MDS follows SDS, where CFG value is set to 100. We stop the training if optimized parameter values stabilize within one order of magnitude. The training process requires around 30 iterations. The iteration time highly depends on the number of input Gaussian kernels, and it is within 30 seconds for most cases.
@@
Figure 3: (a) Text-conditioned optimization; (b) Image-conditioned optimization. Right images are the space-time (X-t) slices, one axis represents time and the other axis shows a space slice (red line) of the object.
@@
Comparison with State-of-the-art Works. We report the quantitative results of all the compared methods in Table 1. Since PhysDreamer hasn’t released its training implementation, we can only compare four evaluation scenes, where the corresponding ground-truth videos are provided in the video demo. Considering that other methods don’t have extra text inputs, we use the first frame as the image condition to conduct the optimization. According to the evaluation of aesthetic quality, our results are the closest to the ground truth. PhysDreamer has a lower score compared with PhysGaussian, which indicates that pre-generated videos may not be a proper ground truth for supervision. The generation quality of DreamGaussian4D is the worst because its deformation prediction didn’t consider physical constraints.